OpenStudy (anonymous):

ODE Decay Question : In the rock being examined there is 100 mg of uranium and 14 mg of lead. Uranium decays with a half-life of 4.5 *10^9 years and after full decay of 238g of uranium 206 g of lead is left. Determine the rock age. Assume that when the rock formed it contained no lead, and neglect all intermediate decay-chain elements( since their half-life is much smaller than uranium )

OpenStudy (anonymous):

the amount of lead formed is proportional to the amount of Uranium left. the amount of uranium left is equal to the original amount minus the amount decayed. ie \[Pb(t)=K*U(t)\] for any t, and \[U(t)= U(0)- U(0)*\left( \frac{ 1 }{ 2} \right)\ ^{-\frac{ t }{ HL }} \] Can you take it from here ?